Grant counted 277 distinct beluga whales in the St. Lawrence Seaway in one week. The following week he counted 343 distinct belugas. What was the percent (round to the nearest one hundredth of a percent) of increase in the whales counted?
Answer by Theo(13342)

counted 277 the first time.

counted 343 the second time.

difference is 343 - 277 = 66

difference divided by 277 = 66 / 277 = .238267148

percent of increase = rate of increase * 100% = 23.8267148

rounded to the nearest 1 hundredth percent = 23.83%